OpenStudy (anonymous):

solve the equation in the set of complex numbers. give only exact answers, not approximations... x^(-2)-x^(-1)=(3/4)

OpenStudy (anonymous):

how would you go about with the negative exponents?

OpenStudy (anonymous):

x^(-n)=1/(x^n)

OpenStudy (anonymous):

then how would you get rid of fractions?

OpenStudy (anonymous):

\[(2\pm 2\sqrt{2}i)/3\]

OpenStudy (anonymous):

how'd you get that..?

OpenStudy (anonymous):

solving the equation... :D

OpenStudy (anonymous):

nice haha but im still confused on what to do after the x's are on the bottom part of the fractions

OpenStudy (anonymous):

make the bottom part x^2 u will get like this: (1-x)/x^2=3/4 after cross multiplication, u can get a quadratic expression, solve that u will get the answer :)

OpenStudy (anonymous):

ok so i got to the (1-x)/x^2=3/4 what do i do next? O.o

OpenStudy (anonymous):

cross multiplication..

OpenStudy (anonymous):

\[x^{-2}-x^{-1}=(3/4)\]\[\Leftrightarrow\frac{1}{x^2}-\frac{1}{x}=\frac{3}{4}\]

OpenStudy (anonymous):

agree with dalvoron. Crossed multiply it and get x-x^2=3/4. then transpose 3/4 to the other side:)

OpenStudy (anonymous):

am i supposed to get \[1\pm \sqrt{2}i/2\]?

OpenStudy (anonymous):

No need for complex numbers here.\[\frac{1}{x^2}-\frac{1}{x}=\frac{3}{4}\]\[\Leftrightarrow \frac{4}{x^2}-\frac{4}{x}=3\]\[\Leftrightarrow 4-4x=3x^2\]

OpenStudy (anonymous):

thanks! i just use the quadratic formula and then get both answers :D

OpenStudy (anonymous):

Quite right, quite right. Regular factorising is possible also. What did you get as your solutions for x?

OpenStudy (anonymous):

2/3,-2

OpenStudy (anonymous):

Correct, good job!

OpenStudy (anonymous):

THANKS :D

OpenStudy (anonymous):

i have a few more problem that i cant get -.-